# Beyond Efficiency: Scaling AI Sustainably

### v1

@@ I From Warehouse Scale Computing to AI
Large-scale computing infrastructures today are extremely efficient [45, 27], with Power-Usage-Effectiveness (PUE) of roughly 1.1. Performance-per-watt energy efficiency of microprocessors are also steadily improving. For GPUs, the theoretical GFLOPS performance-per-watt doubles every 3 to 4 years [56]. Figure 1 illustrates the notable GPU performance improvement (in FP32 GFLOPS) as a combined effect of higher transistor density, higher frequency, and larger die size. Decades of efficiency optimization across the various dimensions of computer systems has led to orders-of-magnitude energy efficiency improvement for computing.
@@ III Efficiency and Beyond
    As the performance of GPUs continues to improve, the overall model execution time will increasingly depend on the effectiveness of the communication collectives and the underlying networking technologies. Figure 4(b) illustrates that, across Deep Learning Recommendation Models (DLRMs), LLMs, and Multi-Modal (MM) AI technologies, communication optimization can also lead to substantial training throughput improvement [35]. However, we must balance the operational efficiency improvement with embodied carbon stemmed from newer generation of hardware to optimize for total lifecycle emissions.
@@
  • Power capacity is the limiting factor at-scale. The rapid growth in computing is putting significant stress on power grids. Power grids, such as EirGrid in Ireland, may not have energy production capacity to source large power loads of datacenter computing [36]. Datacenters already used around 900 MWh of electricity in Ireland, corresponding to 14% of Ireland’s total electricity used in 2021 [15] and the datacenter electricity need is likely increased to 30% of Ireland’s annual electricity supply by 2029 [23]. For other power grids that are able to support the large power capacity requirement of datacenters, making it cost-effective and environmentally-sustainable is a key challenge. In northern Virginia, where the largest concentration of datacenters in the world is at, Dominion Energy – the Virginia Electric and Power Company considered grid infrastructure designs that trade off cost for carbon emissions [20], from building renewable energy and storage for the power grid to extending the life of coal and natural gas energy generation infrastructure while also considering small modular nuclear reactors [10].

### v2

@@ I From Warehouse Scale Computing to AI
Large-scale computing infrastructures today are extremely efficient [45, 27], with Power-Usage-Effectiveness (PUE) of roughly $1.$ Performance-per-watt energy efficiency of microprocessors are also steadily improving. For GPUs, the theoretical GFLOPS performance-per-watt doubles every 3 to 4 years [56]. Figure 1 illustrates the notable GPU performance improvement (in FP32 GFLOPS) as a combined effect of higher transistor density, higher frequency, and larger die size. Decades of efficiency optimization across the various dimensions of computer systems has led to orders-of-magnitude energy efficiency improvement for computing.
@@ III Efficiency and Beyond
    As the performance of GPUs continues to improve, the overall model execution time will increasingly depend on the effectiveness of the communication collectives and the underlying networking technologies. Figure 4(b) illustrates that, across Deep Learning Recommendation Models (DLRMs), LLMs, and Multi-Modal ($M$) AI technologies, communication optimization can also lead to substantial training throughput improvement [35]. However, we must balance the operational efficiency improvement with embodied carbon stemmed from newer generation of hardware to optimize for total lifecycle emissions.
@@
  • Power capacity is the limiting factor at-scale. The rapid growth in computing is putting significant stress on power grids. Power grids, such as EirGrid in Ireland, may not have energy production capacity to source large power loads of datacenter computing [36]. Datacenters already used around 4,000 GWh of electricity in Ireland, corresponding to 14% of Ireland’s total electricity used in 2021 [15] and the datacenter electricity need is likely to increase to 30% of Ireland’s annual electricity supply by 2029 [23]. For other power grids that are able to support the large power capacity requirement of datacenters, making it cost-effective and environmentally-sustainable is a key challenge. In northern Virginia, where the largest concentration of datacenters in the world is at, Dominion Energy – the Virginia Electric and Power Company – considered grid infrastructure designs that balance between cost and carbon emissions [20]. It does so by considering a wide variety of energy generation options, from building renewable energy and storage for the power grid to extending the life of coal and natural gas energy generation infrastructure while also considering small modular nuclear reactors [10].